# Synthesize-on-Graph: Knowledgeable Synthetic Data Generation for Continue Pre-training of Large Language Models

### v2

@@ 4.5.1 Influence Over Different Generation Strategy
(c) Entity distribution: 6M SoG synthetic data with CoT and CC generation.
@@
Dataset CoT + CC CoT CC Direct QA HHRAG(X1.51.51.51.5) $70.9$ $70.6$ $63.7$ $55.3$ HHRAG(X4.54.54.54.5) $74.1$ $72.9$ $62.6$ $55.3$ QuALITY(X1.51.51.51.5) $44.0$ $44.7$ $38.9$ $37.4$ QuALITY(X4.54.54.54.5) $46.2$ $47.5$ $42.8$ $37.4$
Table 2: CPT vs. RAG results: LLM denotes Llama-3-8B-Instruct. CPT denotes the SoG synthetic CPT. Zero-shot denotes direct answering.
@@
4.6 CPT vs. RAG
In this experiment, we aim to answer whether non-parametric external knowledge in retrieval-augmented generation (RAG) can be replaced by parametric knowledge acquired through SoG-based CPT. Specifically, we adopt Llama-3-8B-Instruct as the base model and evaluate its performance on the MHRAG task under three configurations: LLM with SoG CPT, LLM with RAG, and LLM with both SoG CPT and RAG. From the results in Table 2, both RAG and CPT individually bring significant and similar performance gains to the LLM. Interestingly, applying RAG on top of the LLM already enhanced by synthetic CPT does not lead to further improvements. In fact, this combined setting performs worse than using either method alone. We argue that although RAG still holds a marginal advantage in performance, this advantage is outweighed by the broader benefits of synthetic CPT—including eliminating the need for retrieval, enabling shorter input windows for higher efficiency, and saving considerable computational costs in long term (RQ4). Finally, another interesting observation is that on both datasets, models CPT only on the original corpus perform worse than the original model, see Raw CPT in Figure 3. This may be due to the lack of data diversity and distribution differences from the original model’s pretraining data. This further highlights the importance of Synthetic CPT. Based on this discussion, while not yet ready to conclude that RAG can be fully replaced by CPT, our results demonstrate the substantial potential of parametric knowledge.
5 Limitations
While our method shows promising results, several limitations remain. First, although we conducted experimental analysis on the setting of sampling path length in MHRAG, this setting is task-dependent, and determining an appropriate setting for different datasets may require empirical tuning. Second, continue pretraining may introduce unstable LLM output, which requires additional training techniques Ke et al. (2023). We leave these for future work.
6 Conclusion
We propose Synthetic-on-Graph (SoG) framework, a context-graph-enhanced synthetic data generation method that effectively incorporates cross-document knowledge associations, which combine balanced sampling with Chain-of-Thought and Contrastive Clarifying generation strategies. Experimental results show that SoG achieves SOTA performance on multi-hop QA tasks while showing better generalization capability. Our work highlights the potential of SoG as a scalable and efficient solution for continue pretraining, offering new directions for optimizing large language model training in knowledge-intensive domains.
Appendix A Appendix
A.1 Influence of Path Length
We conduct a comparison to assess the impact of different sampling path length choices on the performance of CPT training in Table 3. The 1-hop paths can generate up to 5×5$\times$5 × the data volume; therefore, only the 4.5×4.5$\times$4.5 × result is reported. In general, the 1-hop setting achieves the best performance. The data synthesized from 2-hop paths also show significant performance. However, the 3-hop paths perform considerably weaker. We believe that this may be related to the inherent difficulty of the dataset’s tasks. Furthermore, considering the challenges of constructing multi-hop reasoning tasks, most reasoning tasks are designed within two hops Ma et al. (2025).
@@ A.1 Influence of Path Length
Scale 1-Hop 2-Hop 1+2-Hop ($1:1$) 3-Hop 4.5×4.5$\times$4.5 × $74.0$ $71.9$ $72.5$ $69.3$ 9×9$\times$9 × - $73.5$ $76.1$ $70.7$
A.2 Configuration Adjustment Detail for QuALITY
@@
A.3 Balanced Secondary Sampling
@@
[…]
@@
1:Input: remaining paths set =$\mathcal{R}=\{P\}$, target coverage rate $r$, EntityUtilizationDict, standard length $l$, and entity to chunk index EntityToChunk.
@@
[…]
@@
21:         Add$(\mathcal{P^{*}}[\text{``cot''}][\texttt{cut:}])\text{ back to }\mathcal{R}$
@@
[…]
@@ A.4 Prompt
Figure 6: CC Synthetic Prompt
A.5 Cases

### v3

@@ 4.5.1 Influence Over Different Generation Strategy
(c) Entity distribution: 6M SoG synthetic data with CoT and $C$ generation.
@@
Dataset CoT + CC CoT CC Direct QA MHRAG(X1.51.5) $70.9$ $70.6$ $63.7$ $55.3$ MHRAG(X4.54.5) $74.1$ $72.9$ $62.6$ $55.3$ QuALITY(X1.51.5) $44.0$ $44.7$ $38.9$ $37.4$ QuALITY(X4.54.5) $46.2$ $47.5$ $42.8$ $37.4$
Table 2: CPT vs. RAG results: Base LLM denotes Llama-3-8B-Instruct. CPT LLM denotes the model CPT on the SoG data. Zero-shot denotes directly answering by the corresponding model.
Model RAG Zero-shot Base LLM $73.5$ $55.3$ CPT LLM $70.7$ $73.2$
4.6 CPT vs. RAG
In this experiment, we aim to answer whether non-parametric external knowledge in retrieval-augmented generation (RAG) can be replaced by parametric knowledge acquired through SoG-based CPT. Specifically, we adopt Llama-3-8B-Instruct as the base model and evaluate its performance on the MHRAG task under three configurations: LLM with SoG CPT, LLM with RAG, and LLM with both SoG CPT and RAG. From the results in Table 2, both RAG and CPT individually bring significant and similar performance gains to the LLM. Interestingly, applying RAG on top of the LLM already enhanced by synthetic CPT does not lead to further improvements. In fact, this combined setting performs worse than using either method alone. We argue that although RAG still holds a marginal advantage in performance, this advantage is outweighed by the broader benefits of synthetic CPT—including eliminating the need for retrieval, enabling shorter input windows for higher efficiency, and saving considerable computational costs in long term (RQ4). Our findings highlight that incorporating SoG synthetic data into CPT enables parametric knowledge to streamline task adaptation and enhance output controllability, offering a more efficient alternative to reliance on inference-time retrieval.
5 Conclusion
We propose Synthesize-on-Graph (SoG) framework, a context-graph-enhanced synthetic data generation method that effectively incorporates cross-document knowledge associations, which combine balanced sampling with Chain-of-Thought and Contrastive Clarifying generation strategies. Experimental results show that SoG achieves SOTA performance on multi-hop QA tasks while showing better generalization capability. Our work highlights the potential of SoG as a scalable and efficient solution for continued pretraining, offering new directions for optimizing large language model training in knowledge-intensive domains.
Appendix A Appendix
A.1 Datasets
  • MultiHop-RAG (MHRAG[19] is specifically designed to challenge the multi-hop reasoning capabilities of LLMs. It consists of queries constructed from news articles published between September and December 2023, which include information beyond the training cutoff of existing LLMs, ensuring that synthetic data is required to fill knowledge gaps. In addition, each query requires models to integrate evidence from multiple documents, mimicking real-world scenarios where knowledge is dispersed across sources. Existing LLMs, even RAG systems, often struggle with such tasks, underperforming in tasks that demand integrating and reasoning over scattered evidence. This dataset serves as an ideal benchmark to evaluate how SoG-generated synthetic data equips LLMs to utilize their internal knowledge for handling complex multi-hop reasoning effectively.
  • BioASQ [20]: The BioASQ question answering (QA) benchmark dataset contains questions in English, along with golden standard (reference) answers and related material. The dataset has been designed to reflect real information needs of biomedical experts, assess the comprehensive understanding of professional knowledge, and is therefore more realistic and challenging than most existing datasets. We aim to explore challenging problems in professional domains that require highly specialized expertise, and investigate to what extent SoG can provide models with better learning corpora.
  • QuALITY [21] is a multiple-choice question-answering dataset for long document comprehension. Unlike in prior work with passages, the questions are written and validated by contributors who have read the entire passage, rather than relying on summaries or excerpts. For a fair comparison with the state-of-the-art CPT synthetic data method, EntiGraph, we also chose this dataset for evaluation.
A.2 Performance on More Backbone Models
Evaluating across more base models is crucial for assessing the robustness and generalizability of SoG. To this end, we have conducted additional experiments using Qwen2.5-7B-Instruct and Qwen2.5-32B-Instruct on the MHRAG dataset. The results, presented below, show that, with SoG CPT, smaller models tend to yield closer performance to the larger model:
Table 3: Performance of SoG on the MHRAG dataset across different backbone models.
@@ A.2 Performance on More Backbone Models
Qwen2.5-3B-Instruct 46.7 67.1  (+43.7%) 73.0  (+56.4%)
Qwen3-8B 50.7 70.5  (+39.0%) 76.4  (+50.7%)
LLaMA-3-8B-Instruct 48.7 70.9  (+45.6%) 75.4  (+54.8%)
Qwen2.5-32B-Instruct 55.6 73.4  (+32.0%) 81.3  (+46.2%)
Table 4: Performance of SoG on the BioASQ dataset across different backbone models.
[…]
Qwen2.5-3B-Instruct 11.8 21.7  (+83.9%) 29.4  (+149.2%)
Qwen3-8B 10.3 20.9  (+103.9%) 28.5  (+176.7%)
LLaMA-3-8B-Instruct 13.5 26.2  (+94.1%) 35.1  (+159.3%)
Qwen2.5-32B-Instruct 27.8 44.5  (+60.1%) 57.3  (+106.1%)
A.3 Influence of Path Length
We conduct a comparison to assess the impact of different sampling path length choices on the performance of CPT training in Table 5. The 1-hop paths can generate up to $5\times$ the data volume; therefore, only the $4.5\times$ result is reported. In general, the 1-hop setting achieves the best performance. The data synthesized from 2-hop paths also show significant performance. However, the 3-hop paths perform considerably weaker. We believe that this may be related to the inherent difficulty of the dataset’s tasks. Furthermore, considering the challenges of constructing multi-hop reasoning tasks, most reasoning tasks are designed within two hops [22].
@@ A.3 Influence of Path Length
Scale 1-Hop 2-Hop 1+2-Hop ($1:1$) 3-Hop $4.5\times$ $74.0$ $71.9$ $72.5$ $69.3$ $9\times$ - $73.5$ $76.1$ $70.7$
A.4 Configuration Adjustment Detail for QuALITY
@@
A.5 Implementation Cost
Our method does not rely on the strongest or most expensive LLMs. All generations are conducted with GPT-4o-mini, a fast and cost-efficient model (pricing: \$0.15 per 1M input tokens, \$0.08 per 1M cached input tokens, and \$0.60 per 1M output tokens). In the synthetic generation stage, the average input and output token counts per instance are approximately 1,7001{,}700 and $900$, respectively. Based on our experiments, expanding the corpus by $3\times$$4.5\times$ (i.e., $\approx 2$33M tokens for Enti-Graph) is already sufficient to yield substantial performance improvements. Consequently, the overall cost of SoG remains modest, making it a practical and accessible choice even under limited computational or financial resources.
A.6 Limitations
While our method shows promising results, several limitations remain. First, although we conducted experimental analysis on the setting of sampling path length in MHRAG, this setting is task-dependent, and determining an appropriate setting for different datasets may require empirical tuning. Second, continued pretraining may introduce unstable LLM output, which requires additional training techniques [23]. We leave these for future work.
A.7 Balanced Secondary Sampling
@@
7:       $\texttt{BalancedSampling}(\mathcal{R},r,\texttt{EntityUtilizationDict},l,\texttt{EntityToChunk})$
@@
1:Input: remaining paths set $\mathcal{R}=\{P\}$, target coverage rate $r$, EntityUtilizationDict, standard length $l$, and entity to chunk index EntityToChunk.
@@
7:       $\textsc{PathUtilizationCount}(P)=\sum_{\texttt{node}\in P}\texttt{EntityUtilizationDict}[\texttt{node}])$
8:  # In default, $l=\texttt{TotalNumberOfChunksInCorpus}/(\texttt{hop}+1)$ and $r=100\%$
@@
22:   Reverse EntityUtilizationDict based on $\mathcal{P^{*}}[\text{``cot''}][\texttt{cut:}]$
23:   $\mathcal{P^{*}}[\text{``cot''}]\leftarrow\mathcal{P^{*}}[\text{``cot''}][0\texttt{:cut}]$
24:   $\texttt{SparseEntities}\leftarrow\texttt{EntityUtilizationDict}[0\texttt{:}k]$
@@ A.8 Prompt
Figure 6: $C$ Synthetic Prompt
A.9 Cases